# Understanding long-term energy use in off-grid solar home systems in sub-Saharan Africa

### v1

@@ Abstract
Solar home systems provide low-cost electricity access for rural off-grid communities. As access to them increases, more long-term data becomes available on how these systems are used throughout their lifetime. This work analyses a dataset of 1,000 systems across sub-Saharan Africa. Dynamic time warping clustering was applied to the load demand data from the systems, identifying five distinct archetypal daily load profiles and their occurrence across the dataset. Temporal analysis reveals a general decline in daily energy consumption over time, with 57% of households reducing their usage after the first year of ownership. On average, there is a 33% decrease in daily consumption by the end of the second year compared to the peak demand, which occurs on the 96th day. Combining the load demand analysis with payment data shows that this decrease in energy consumption is observed even in households that are not experiencing economic hardship, indicating there are reasons beyond financial constraints for decreasing energy use once energy access is obtained.
@@ II Literature Review
In addition to understanding day-to-day energy use, it is important to consider long-term behaviour, particularly in the context of energy access for previously unelectrified households. In these situations, behaviour is more unpredictable [36, 37]. Therefore, long-term load demand analysis is imperative to gain insights into how systems deployed in LMICs are used. It is commonly thought that energy demands will increase as access is granted [36, 37, 38], however, this conjecture needs to be supported by real-world data [39]. Bisaga et al. [40] compared energy use of households across Rwanda, splitting them into groups depending on the length of time they had been using systems. They found that households that had systems for over a year generally used less energy each day than those who had obtained systems within the last 6 months, despite the more established households having generally more appliances. Kizilcec et al. [41] also conducted work on Rwandan SHS energy use, finding a reduction in electricity demand over the first year of use for the majority of households in their dataset. The decrease in electricity consumption was even more pronounced in households that owned a TV.
@@
Beyond the first year of ownership, Dominguez et al. [42] found a general increase in monthly electricity expenditure once access is gained, along with an increase in appliance ownership, but this plateaued at the 2-3 year mark. Conversely, [43] found that rural households in Rwanda with grid connections decreased their electricity consumption over the first ten years of connection. This analysis includes monthly electrical consumption data on 147,074 rural households across the country, focusing on 174 households for which this data is backed up with household surveys. The surveys found that household appliances largely consist of lighting and entertainment devices, but economically productive appliances are a rarity. Louie et al. [44] analysed the load demand of off-grid households in Navajo Nation over 2 years. They found a large variation both between different households and within households, with seasonal changes having a large impact on intra-household variations. Additionally, energy use in the second year was on average 10.3% lower than in the first year, highlighting longer-term temporal load shifts.
Considering these long term temporal changes in energy use described in literature, the picture is not clear. Increasing consumption is anticipated, along with the increase of appliances and productive uses. But, there are also datasets indicating the opposite.
Although electricity consumption data can indicate trends, further information is required to understand underlying causes. Often the barrier to rural electrification is economic [45, 46]. Mergulhao et al. [47] clustered the payment patterns for SHS users in Rwanda and Kenya, highlighting the diversity of PAYG payment methods. However, there is a gap in linking payment data to energy use and changes over time. Guajardo et al. [48], established a link between energy and payments in SHS, finding decreased probability of late payments for higher energy use customer. However, authors focused on payment data rather than load demand profiles and how they change over time. Lukuyu et al. [35] further examined this link between east African SHS electricity consumption and payments and found that customers with greater monthly payments had greater peak electricity consumption. However, crucially, they did not investigate how households change their spending and usage patterns over time.
@@
From the literature, there are clear gaps this work aims to address. Our contributions are as follows: firstly, there is a need for large-scale analysis of daily SHS load demand profiles to elucidate patterns in real-world data; secondly, we consider the progression of energy use over several years, rather than observing a single snapshot in time; finally, we combine payment and energy data from households to gain insight into the role of economic factors. This research also provides access to the measured data for others to use, supporting the collective efforts of researchers and industry to achieve SDG7.
III Energy use dataset
Data analysed in this paper was provided by BBOXX, and consists of time series from 1,000 SHSs across Africa. The BBOXX SHSs consist of a 50 W solar panel, 12 V 17 Ah lead-acid battery and various loads, including lighting, phone charging, fans, radios and TVs.
The BBOXX dataset contains data for over half a million of these systems, so the 1,000 used for this analysis were chosen by taking a stratified sample from the whole dataset, based on system lifetime to ensure a range of users, old and new, are represented. All SHSs younger than 1 year were excluded so longer term analysis can be done for all customers, and the oldest system in the dataset is 1230 days. The geographical distribution of the systems across Africa is shown in Fig. 1. All the systems are located in sub-Saharan Africa, with most located in Kenya, Togo, Democratic Republic of the Congo and Rwanda.
@@ III Energy use dataset
Figure 1: Geographical distribution of the downsampled SHS used in the clustering process. The countries coloured white had no SHS in the dataset.
@@
While BBOXX collect lots of data from their SHSs, for load demand analysis, the time, output current and voltage are the only required time series. The raw data was measured with non-uniform sampling, whereby a measurement is recorded whenever a change is detected, with the associated time stamp of this change also recorded. This is to second resolution, with a 10-minute backstop if no change is detected in that time. To convert this to hourly load demand data, the following method was used:
@@
    Find instantaneous power $P(t)$ at each time stamp $t$ from the measured voltage $V(t)$ and discharge current $I_{\text{out}}(t)$,
    $P(t)=V(t)\times I_{\text{out}}(t)$ (1)
@@
    Determine the energy consumed by integrating the power. For hourly intervals, $\Delta t=1$ hour,
@@
    Each household is then represented by a discrete energy vector, $E$, where each element is the energy consumed in the respective hour. $t_{end}$ represents the final time step in the time series data.
    $E_{k}=E(k\cdot\Delta t)\quad\text{where}\quad k\in\left[1,\frac{t_{\text{end}}}{\Delta t}\right]$ (3)
For each of the 1,000 SHS, a stratified sample based on daily energy consumption was completed to get 10 days representative from each system. From these 10,010 time series, a further energy consumption based stratified sample was taken to get 2,000 time series. This further stratified sampling step was taken because there is not so much dissimilarity in the daily load profiles that it is beneficial to have evermore time series in the clustering problem. Additionally, it is still feasible to cluster using MIP for a dataset of 2,000 which can confirm if a globally optimal solution is found to the clustering problem. Therefore, the clustering exercise was performed on 2,000 time series, each containing 24 values representing hourly load demand for a 24-hour period, starting at midnight.
Once the clustering problem has been solved, the representative centroids are extrapolated to the full dataset, which contains all days for the 1,000 SHSs. To do this, each household is broken down into their daily time series, and each of these daily time series are clustered, using DTW, into the clusters previously found. This gives a chronological sequence of days in each cluster, from the acquisition of the SHS to the time of analysis.
The BBOXX dataset also contains payment data for their customers, extensively analysed in [47]. For this work, the important payment information is the time series of the remaining credit on each SHS. The contracts vary both by location and households’ appliances - households’ energy limits are restricted based on the number of appliances they own, which relates to their tariff. To allow comparison between households, we convert the credit in local currency to the number of days of electricity use remaining on the system. This removes the need for conversion between countries currencies and the different magnitude of payments. Once a household runs out of credit, if it is not topped up again, the system goes into ‘disabled’ state where discharging is not permitted. When a SHS enters this state, it is deemed an economic outage [49].
Finally, we obtain one more metric from the payment data; the utilisation rate. This is defined as the number of days a household is not in economic outage divided by the number of days the household has owned the system. For example, a household with a utilisation rate of 0.9 will be in economic outage on 10% of days since obtaining the SHS. Each household has a utilisation rate calculated for it, at the time of analysis.
@@ IV Clustering Method
This work uses a DTW-MIP clustering software, as described in [50]. The details of the problem formulation can be found in the original paper, but below we present the key ideas here for readability and completeness.
The DTW distance between 2 time series, $x$ and $y$ of length $n$ and $m$, respectively, can be found using dynamic programming:
@@
The final element $c_{n,m}$ is the total DTW distance between $x$ and $y$. In this work, we chose to not use a warping window, which restricts which parts of one time series can be mapped to another [25]. This is because of the benefits of allowing warping along the whole day, allowing similarity to be found between a peak whether it is at night or in the morning. DTW has an implicit penalty for warping, as more warping between two signals means more pairwise comparisons embedded in the final cost. Therefore, while there is an implicit warping penalty, the hard constraint of a warping window is not desirable.
The DTW distance between each time series in the clustering problem is stored in a symmetric matrix $D_{p\times p}$ where $p$ is the number of time series in the clustering problem. A binary square matrix $A_{p\times p}$ is used for the MIP cluster assignment, where $A_{ij}=1$ if time series $j$ is in the cluster with centroid $i$. The clustering optimisation can be solved by
$A^{\star}=\arg\min_{A}\sum_{i}\sum_{j}D_{ij}\times A_{ij}$ (5)
To choose the value of $k$ (i.e. the number of clusters), the clustering exercise was done for 2 to 8 clusters. We then compared the costs and silhouette scores [51] to decide on a $k$ value of 5.
@@ V Results and Discussion
The results of this work are divided into three sections. The first section presents representative daily load demand profiles identified through clustering. The second section focuses on the long-term changes in households energy use. This is done in two subsections; the first looks at consumption trends, while the second applies the results of the clustering to the entire dataset, resulting in each household being represented by a series of cluster labels. We examine how each household moves within the cluster space over time, from the initiation of ownership to the time of analysis. The third section considers the payment data and its links to electricity use.
@@ V-A Daily clustering
Fig. 2 depicts the centroids of the five clusters resulting from the DTW-MIP clustering and the heatmaps for each cluster. The centroids represent the individual time series that minimise the distance between the clusters, as described in the methodology section. Consequently, they can be regarded as the representative time series for the cluster. Nevertheless, it is crucial to examine all the time series within each cluster in order to gain further insight into the discrepancies between the clusters. The heatmaps present each time series within each cluster as a row, with lighter blues representing low energy consumption and darker blues indicating higher energy consumption, which confirms the general agreement between the cluster representatives and their time series.
Table I shows the division of time series in the subset between the clusters, along with the results from the extrapolation of the centres from the subset to the whole dataset. The validity of the extrapolation is shown here, with a larger proportion of time series in the economic outage subset of low use, but generally there is a consistent representation of the full dataset in the subset.
@@
TABLE I: Comparison of the clustering results from the subset, to when the centres are extrapolated to the full dataset.
@@
TABLE II: Mean characteristics of all time series in the dataset, separated into clusters.
Table II shows important characteristics of the time series within each cluster, analysed from the whole dataset, including the peaks in the time series, and information on appliances and time of energy use. A peak within a time series is defined as having double the amplitude of the mean energy consumption of the time series, at least 2 hours apart from another distinct peak (this is to reduce the risk of fluctuations across the threshold counting as two peaks). Additionally, to prevent users with a very small consumption returning lots of peaks due to a very low mean consumption, a peak must have a minimum amplitude of 10W. Defining the cut-off for what is and is not a peak is not rigid, and therefore it is important to also consider the heatmaps in Fig. 2, where the additional peak in the double peak cluster is evident.
@@
Analysing the individualities of each cluster, low use stands out. Given the operational conditions of PAYG systems, this cluster signals two potential scenarios: either there is an economic outage, or there is virtually no usage despite payment. To differentiate between these two states, a new cluster was created, classifying low use time series as economic outage if the time credit balance was 0. It is notable that 24% of days in the dataset belong to economic outage and low use combined, with 75% of them corresponding to economic outages.
@@
The moderate use cluster is a large cluster and has the next lowest mean daily consumption. Looking at the heatmap for this cluster in Fig. 2, the typical use is characterised by consistent low use, with more energy consumed in the evening for most households and a slight increase in use during the morning too. The evening use spans between approximately 3pm to 9pm and therefore crossed over the day/night barrier of 6pm, leading to a non-distinct preference towards daytime or nighttime use. The peak demand is fairly low, and few households within this cluster have TVs, which is reflected in the low mean appliance power. It can be assumed that households with TVs and higher appliance power are not using all their appliances on days when they are in the moderate use cluster.
Nighttime use is another large cluster, with a slightly larger mean daily consumption than moderate use, however there is significant overlap in the daily energy consumption between these two clusters. Nighttime use rate of households TV ownership is low, much like in moderate use, and have a low appliance power of 55 W. The characteristic that distinguishes a time series from being in moderate use or nighttime use, is that nighttime use profiles consume most of their energy at night – between the hours of 6pm and 6am. DTW, as a shape based method, is a very effective measure for picking up discrepancies like this as the shape changes between a generally daytime use profile and nighttime use profile. A nighttime profile will have higher use at the start and end of the time series and lower consumption in the middle, creating a U shape. Whereas the daytime use profile will have lower use at the start and end of the time series with more use somewhere in the middle, creating more of an upside down U shape. The benefit of DTW is the time of peaks, or the beginning/end of the increase or decrease in energy consumption, does not have to align exactly.
@@
The final two clusters are higher consumption days. The mean consumption is marginally higher in double peak, and both clusters have similar rated power appliances. Households in single peak and especially in double peak are very likely to have a TV, which is the main high wattage appliance. However, the difference between these two clusters is the number of peaks. Time series in double peak generally have 2 peaks, while most time series in single peak have only 1. This is seen both in the mean daily peaks in Table II and the heatmaps in Fig. 2. A final distinction between these two higher consumption clusters is that there is a higher proportion of daytime use in double peak, caused by the additional morning peak alongside the evening peak.
@@
Finally, we include the proportion of days in each cluster that are weekdays in Table II to show there is no significant correlation – for each cluster the weekday proportion is around 72% which is the same as the proportion of weekdays to weekend days. This differs to HICs where there is often a noticeable difference in weekday and weekend consumption [52].
Considering all the clusters together, there is a significant non-homogeneity between households energy use, despite the equal technical capabilities of each system with same size battery and solar panel. It can be speculated that the causes of this large variation are financial, or due to the diverse energy needs of different households on different days. However, equipped with this knowledge of the large load demand variation, both on an inter- and intra-household level, it follows that having the same technical specification for each system may be illogical.
@@ V-B1 General consumption trend
Figure 3: The mean daily household consumption across the SHS initially increases when households first acquire their SHSs, but over time, the consumption decreases. The confidence interval decreases overtime as the number of households in the sample decreases. The bottom subplots split the households into high energy ($>50$ Wh of appliances) and low energy ($\leq 50$ Wh of appliances).
Fig. 3 shows an initial increase in the mean daily consumption, peaking on the 96th day. This initial increase is likely a ‘settling in period’ while the system is getting set up, as well as households purchasing appliances and becoming familiar with the system.
In the first 6 months, daily consumption generally increases for 58% of customers, but by the end of the first year, only 43% of customers continue to increase their energy consumption. Beyond the first year, 77% of users have generally decreasing daily energy consumption since obtaining the system up until the point of data retrieval (between 365 and 1200 days). In absolute terms, the mean energy consumption over two years shows a 33% reduction from its peak, dropping from 66 Wh/day to 44 Wh/day. This is contrary to common thought in literature, which expects energy consumption to increase once access is granted, generally thought to be because users will acquire more appliances and use the system increasing amounts, therefore increasing energy consumption [42]. However, this is not seen here. Instead, SHS users appear to follow a long-term consumption trend similar to that witnessed in their grid-connected counterparts [43].
In the top graph in Fig. 3 we see the general decrease in consumption over time. This is particularly prominent in the first 400 days. Beyond this point, the trend diverges depending on the household. Households are split into high and low energy, based on the summation of all household appliance power. High energy households have appliances that sum to over 50 Wh, while low energy households are equal to or below this.
@@
The high energy households’ mean consumption begins to plateau beyond the 400-day mark, when considering all days. When economic outages are removed from the calculation, there is still a slight decrease in consumption, especially beyond 700 days. Therefore, the stabilisation of the mean consumption is assisted by a decrease in the impact of economic outages beyond this point in high energy households, as well a more consistent consumption when the systems are in use. Another important observation from the high energy subplot is that when economic outages are ignored, the consumption continues to increase until around 250-days of ownership. The decrease in the mean consumption before this time shows a high proportion of economic outages, so when the system is in a use, the consumption increases but the significance of economic outages is so large here that the total mean actually decreases.
@@
In low energy households, the mean relative decrease until the end of the first year is even greater (24%) than in high energy households (12%). The decrease continues, although at a lower rate, until around 800-days. The difference between the mean consumption with and without economic outages in low energy households grows quickly in the first part of ownership and then gradually increases over time.
@@
A decrease in energy use overtime can lead to oversized systems. This is undesirable for many reasons – the solar potential of the system isn’t realised, so there is wasted energy [53, 54], the batteries are not cycling in an optimal manner which can shorten their lifetime [55, 56] and customers are paying for a larger system than they need, increasing the cost of their energy. This is particularly detrimental when the cost of energy is a significant barrier to electrification [57, 46].
V-B2 Cluster space trend
Beyond increasing and decreasing trends, the long-term use for individual households can be analysed by finding an individual households movements within the cluster space over time. From this analysis, an interesting insight is the lack of homogeneity within households. Homogeneity here is defined as the fraction of days a household’s load demand profile is in its dominant cluster (the dominant cluster being the cluster a household is in for the highest number of days). Table III shows the characteristics of households, split into their dominant cluster, including the homogeneity. The mean homogeneity across all households is 55%, so 45% of the days, households aren’t using energy how they ‘typically’ would.

### v2

@@ Abstract
Solar home systems provide low-cost electricity access for rural off-grid communities. As access to them increases, more long-term data becomes available on how these systems are used throughout their lifetime. This work analyses a dataset of 1,000 systems across sub-Saharan Africa. Dynamic time warping clustering was applied to the load demand data from the systems, identifying five distinct archetypal daily load profiles and their occurrence across the dataset. Temporal analysis reveals a general decline in daily energy consumption over time, with 77% of households reducing their usage compared to the start of ownership. On average, there is a 33% decrease in daily consumption by the end of the second year compared to the peak demand, which occurs on the 96th day. Combining the load demand analysis with payment data shows that this decrease in energy consumption is observed even in households that are not experiencing economic hardship, indicating there are reasons beyond financial constraints for decreasing energy use once energy access is obtained.
@@ II Literature Review
In addition to understanding day-to-day energy use, it is important to consider long-term behaviour, particularly in the context of energy access for previously unelectrified households. In these situations, behaviour is more unpredictable [36, 37]. Therefore, long-term load demand analysis is imperative to gain insights into how systems deployed in LMICs are used. It is commonly thought that energy demands will increase as energy access is obtained [36, 37, 38], however, this conjecture needs to be supported by real-world data [39]. Bisaga et al. [40] compared energy use of households across Rwanda, splitting them into groups depending on the length of time they had been using systems. They found that households that had systems for over a year generally used less energy each day than those who had obtained systems within the last 6 months, despite the more established households having generally more appliances. Kizilcec et al. [41] also conducted work on Rwandan SHS energy use, finding a reduction in electricity demand over the first year of use for the majority of households in their dataset. The decrease in electricity consumption was even more pronounced in households that owned a TV.
@@
Beyond the first year of ownership, Dominguez et al. [42] found a general increase in monthly electricity expenditure after initial adoption, along with an increase in appliance ownership, but this plateaued at the 2-3 year mark. Conversely, [43] found that rural households in Rwanda with grid connections decreased their electricity consumption over the first ten years of connection. This analysis includes monthly electrical consumption data from 147,074 rural households, focusing on 174 households with survey data. The surveys found that household appliances largely consisted of lighting and entertainment devices, but economically productive appliances were a rarity. Louie et al. [44] analysed the load demand of off-grid households in the Navajo Nation over 2 years and found a large variation both between and within households, the latter being impacted by seasonal changes. Additionally, energy use in the second year was on average 10% lower than in the first year, highlighting longer-term temporal load shifts.
The overall picture for long-term temporal changes in energy use remains unclear—increased consumption is anticipated but there are also datasets indicating the opposite. Further information is required to understand underlying causes. The barrier to rural electrification is often economic [45, 46], but there is a research gap in understanding the link between diverse PAYG payment data [47] and energy use. Existing studies establish a link between energy use and payment, but focus mainly on payment behaviour impacts [48] and very coarse-grained energy consumption data [35].
@@
This work makes several contributions: firstly, a large-scale analysis of daily SHS load-demand profiles is undertaken, elucidating patterns in real-world data; secondly, progression of energy use over several years is considered, rather than observing a single snapshot in time; finally, payment and energy data are combined, enabling insight into the role of economic factors. This research also provides access to the measured data for others to use, supporting the collective efforts of researchers and industry to achieve SDG7.
III Energy use dataset and analysis
[…]
@@ III Energy use dataset and analysis
Figure 1: Geographical distribution of the downselected solar home systems used in the clustering process. The countries coloured white hosted no systems in our dataset.
@@
The time, output current, and output voltage were used for load-demand analysis, with non-uniform sampling, i.e., measurements were recorded whenever a change above a certain threshold was detected. This was at second resolution, with a 10-minute backstop recording taken if no change was detected within that time. The following method was used to convert these measurements to hourly load demands:
@@
    Calculate instantaneous power $P(t)$ at each time $t$ from measured voltage $V(t)$ and load current $I_{\text{out}}(t)$,
    $P(t)=V(t)\times I_{\text{out}}(t).$ (1)
@@
    Determine the energy consumed in one hour by integrating the power over hourly intervals, $\Delta t=1$ hour,
@@
    For each household, construct a vector of hourly energy consumption values, $E$, where each element is the energy consumed in that hour ($t_{end}$ represents the final time step in the time series data), i.e.,
    $E_{k}=E(k\cdot\Delta t)\quad\text{where}\quad k\in\left[1,\frac{t_{\text{end}}}{\Delta t}\right].$ (3)
For the 1,000 systems considered, there are a total of 647,021 daily load profiles. Since many of these are very similar, we downsampled a subset from the total set to improve computational efficiency. Stratified sampling was used, dividing the dataset into subgroups based on daily energy consumption. First, the distribution of daily energy consumption across the entire dataset was found, then time series were randomly selected so as to keep the same distribution of daily energy consumption in the sampled subset. The process was applied twice: Initially, for each household, ten representative days were selected, giving 10,000 time series in total. This ensures that every household and the overall spread of energy consumption are represented. Then, we applied the same method to reduce this set down to 2,000 time series in total. These steps together ensure both a representatively diverse dataset and the option to apply an algorithm to find a global optimum in the clustering process (see next section).
After the cluster centres were found, each of the 647,021 daily load profiles were assigned to their nearest cluster according to DTW distance. The end result is that each household can then represented by a chronological sequence of days, where each day is labelled with the cluster that is the closest match to the load profile of that day.
The BBOXX dataset also contains customer payment data that has previously been analysed extensively [47]. For this work, the relevant payment information is the time series of remaining credit for each SHS. The contracts vary both by location and load—household energy limits are restricted based on the number of appliances owned, which relates to the customer tariff. To allow comparison between households, we convert the credit in local currency to the number of days of electricity use remaining. This removes the need for conversion between currencies and payment amounts. Once a household runs out of credit a system enters a ‘disabled’ state where discharging is not permitted until further payment—this is deemed an economic outage [49]. The clustering exercise previously described ignored this condition (i.e., clustered systems may or may not be in this stage), but we later consider it when analysing the results.
We also calculated another payment metric—the utilisation rate. This is defined as the number of days a household is not in the economic outage state divided by the number of days the household has owned the system. For example, a household with a utilisation rate of 0.9 experienced economic outages on 10% of total days since obtaining the SHS.
@@ IV Clustering Method
This work uses our DTW-MIP clustering algorithm, fully described in Kumtepeli et al. [50]. A brief overview of the approach is given here for readability and completeness.
The DTW distance (cost) between individual elements of two time series $x$ and $y$ of length $n$ and $m$ (indexed by $i$ and $j$), respectively, is given by
@@
where the final element $c_{n,m}$ is the total DTW distance between series $x$ and $y$. This problem can be solved using dynamic programming. Here we chose not to use a warping window to restrict the parts of one time series that can be mapped to another [25]. This allows long-range similarities to be found between peaks in energy consumption, i.e., whether they occur at night or in the morning. Dynamic time warping already penalises warping because more warping means more pairwise comparisons that are embedded in the final cost. Therefore the hard constraint of a window is not necessary.
The DTW distance between each time series is stored in a symmetric matrix $D_{p\times p}$, where $p$ is the total number of time series. A binary square matrix $A_{p\times p}$ is used for MIP cluster assignment, where $A_{ij}=1$ if time series $j$ is in cluster with centroid $i$. There are $k$ clusters (5b), and each time series must be in one cluster only (5c). Clusters can be found by solving the optimisation problem
$\min_{A}$ $\sum_{i}\sum_{j}D_{ij}\times A_{ij}$ (5a)
subject $\text{subject to}\quad\sum_{i=1}^{p}A_{ii}$ $=k,$ (5b)
$\sum_{i=1}^{p}A_{ij}$ $=1$ $\forall j\in[1,p],$ (5c)
$A_{ij}$ $\leq A_{ii}$ $\forall i,j\in[1,p].$ (5d)
To choose the value of $k$, the algorithm was repeated for 2 to 8 clusters. We then compared the costs and silhouette scores [51] to select a final value $k=5$.
@@ V Results and Discussion
The results of this study are divided into three sections. The first presents representative daily load demand profiles identified through clustering. The second focuses on long-term changes in household energy use, first exploring consumption trends and then long-term trends in cluster allocation over the entire dataset. Finally, we consider links between payment data and electricity use.
@@ V-A Daily clustering
TABLE I: Comparison of clustering results from subset vs. extrapolation to full dataset.
@@
TABLE II: Mean characteristics of all clustered time series.
Fig. 2 depicts the centroids of the five clusters resulting from the DTW-MIP clustering exercise, and heatmaps for each cluster. The centroids are the individual time series at the median of each cluster. It is also helpful to examine all the time series within each cluster to gain further insight into discrepancies between and within clusters. The heatmaps present each time series in each cluster as a row, with lighter blues representing low energy consumption and darker blues indicating high consumption. Overall, the results confirm the general agreement between the cluster representatives and their associated time series.
Table I shows the allocation of time series across clusters within the subset, along with the results from the extrapolation to the whole dataset. The validity of the extrapolation is generally acceptable, with a larger proportion of time series in the economic outage subset of low use, but there is a reasonably consistency between the full dataset and subset.
[…]
@@
The low use cluster stands out as distinct from the other clusters, and signals that either there is an economic outage, or there is virtually no usage despite payment. To differentiate between these two scenarios, a new subset was created, classifying low use time series as economic outage if the credit balance was also zero. It is notable that 24% of days in the dataset belong to economic outage and low use combined. Of these, 75% (i.e., 18%/24%) correspond to economic outages, but there is still a significant proportion of days where the systems are paid for, but not used.
@@
The moderate use cluster is large and has the next lowest mean daily consumption. Typical moderate use (heatmap, Fig. 2) is characterised by consistent low use throughout the day, with more energy consumed in the evening and a slight increase in morning use. The evening use spans between 15:00 and 21:00 and therefore crosses sunset at around 18:00, but drops to almost no consumption after midnight. The peak demand is fairly low and few households with days in this cluster have TVs, which is reflected in the low mean appliance power. Alternatively, households with TVs and higher power appliances are not using all them on days when they are in this cluster.
[…]
@@
The final two clusters, double peak and single peak, represent higher consumption (households are likely to have a TV), with the double peak case slightly higher than the single. And, of course, the number of peaks is different between the two clusters—as seen in Table II and Fig. 2. A final distinction between these two higher consumption clusters is that there is a higher proportion of daytime use in double peak, caused by the additional morning peak alongside the evening peak.
@@
Finally, we include the proportion of days in each cluster that are weekdays (Table II), showing that there is no significant difference between clusters—in all cases the weekday proportion is around 72%, i.e., 5/7. This differs from HICs where there is often a noticeable difference between weekday and weekend consumption [52].
In summary, there is significant heterogeneity in household energy use despite the identical technical capabilities of each system (e.g., same sized battery and solar panel). The causes of this might be financial or due to the diverse energy needs of different households on different days, caused by diversity in activities. There could be a case for supplying different intensity users with differently sized systems.
@@ V-B1 General consumption trend
Figure 3: Top: Mean daily household electricity consumption initially increases, but then decreases. The confidence interval becomes wider over time as the number of households in the sample decreases. Bottom: High-use households ($>50$ Wh of appliances), left, vs. low-use ($\leq 50$ Wh of appliances), right.
[…]
@@
Mean consumption in high-use households begins to plateau beyond 400 days. When economic outages are removed the decrease in consumption is slightly stronger, especially beyond 700 days. Another important observation is that when economic outages are ignored, the consumption in high-use households continues to increase until around 250 days, i.e., the apparent decrease in the mean consumption of high-use households after the peak around day 96 is actually caused by a high proportion of economic outages.
@@
In low-use households, the mean relative decrease to the end of the first year is even greater (24%) than in high-use households (12%). The decrease continues, although at a lower rate, until around 800 days. The difference between the mean consumption with and without economic outages in low-use households grows quickly in the first part of ownership and then gradually increases over time.
@@
A gradual decrease in electricity use can lead to oversized systems. This is undesirable because the full solar generation potential is not realised [54, 55], batteries are not cycled optimally, shortening lifetime [56, 57], and customers pay for a larger system than needed, increasing their electricity cost. This is particularly detrimental because the cost of energy is a significant barrier to electrification [58, 46].
V-B2 Cluster allocation trend
We now analyse long-term trends in cluster allocation for individual households over time. An interesting finding is a lack of homogeneity within households, where homogeneity is defined as the fraction of days that a household’s demand is in its dominant cluster (the cluster the household spends the most time in). Table III shows that the mean homogeneity for all households is only 55%.